OpenStudy (k8lyn911):

Simplify 6/n(18/n((n(n+1)/2) Answer: 54(n+1/n) How exactly do I go about getting that answer?

OpenStudy (k8lyn911):

Equation \[\frac{ 6 }{ n } (\frac{ 18 }{ n } (\frac{ n(n+1) }{ 2 }))\] Answer \[54 (\frac{ n+1 }{ n })\]

OpenStudy (anonymous):